OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!! joey casts a 10 m shadow while his 2 m tall father casts a 25 m shadow. how tall is joey?

OpenStudy (btaylor):

use ratios: \[\frac{x}{10}=\frac{2}{25}\] Cross multiply to get 20 = 25x x = 20/25 = 4/5 m tall